OpenStudy (anonymous):

Farmer Bob has pigs and chickens. He has 37 animals, and there are 124 legs among them altogether. How many chickens does Bob have?

hero (hero):

p + c = 37 4p + 2c = 124

OpenStudy (anonymous):

ok he has

OpenStudy (anonymous):

an also be answered by setting up a system of equations instead of guessing. P + C = 37 (P pigs + C chickens = 37 animals) 4P + 2C = 124 (4P pig legs + 2C chicken legs = 124 legs) Multiply the first equation by -4 and add to the second equation and solve for P.

OpenStudy (anonymous):

A. 25 B. 4 C. 12 D. –4

OpenStudy (anonymous):

if that helps. /.\

OpenStudy (anonymous):

did I help

OpenStudy (anonymous):

yess

hero (hero):

You have to solve for c, not p

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok sooo? is it 25

OpenStudy (anonymous):

I say p 4 example

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

so which one? D;

hero (hero):

You're still confused @wickeddd

OpenStudy (anonymous):

yup

hero (hero):

You must have guessed the right answer instead of solving for it.

OpenStudy (anonymous):

lol yeaah,

OpenStudy (anonymous):

I knew it @wickeddd

OpenStudy (anonymous):

well no cause its the only odd number /.\

OpenStudy (anonymous):

ok?

hero (hero):

Multiply the first equation by 4, you get: 4p + 4c = 148 4p + 2c = 124 Subtract the 2nd from the 1st to get: 2c = 24 c = 24/2 c = 12

hero (hero):

c = number of chickens btw

OpenStudy (anonymous):

im over it thanks for the help.

OpenStudy (anonymous):

u welcome